OpenStudy (anonymous):

wat is the square root of 69 cause i been trying 2figure it out

OpenStudy (anonymous):

idk

OpenStudy (radar):

\[\sqrt{23}\times \sqrt{3}\]

OpenStudy (anonymous):

Very clever. ._.